MIDC/Data Quality Statements
Total Records: 40, displaying 26-40
[Prev Records]
R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
26BMSPhotometer TWC05/17/201114:20:0005/17/201114:48:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
27BMSPhotometer TWC05/07/201105:55:0005/09/201107:58:00There was a planned power outage for SRRL construction activities that started at 5:18 MST. Unfortunately, the backup generator provided to SRRL was providing bad power to the UPS powering the Sci-Tec trackers, therefore the UPS failed at 5:55 MST. Unfortunately, when normal AC power was restored at 11:45 MST, the Sci-Tec trackers had lost all their configuration due to UPS cycling on and off during the outage, therefore were not tracking. The tracker settings were restored on May 9 at 7:58 MST. The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4
28BMSPhotometer TWC05/07/201105:18:0005/07/201111:45:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
29BMSPhotometer TWC04/15/201106:11:0004/15/201106:50:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
30BMSPhotometer TWC03/18/201100:00:0003/18/201107:57:00The Sci-Tec #2 tracker was jammed on the shading plate. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201, TWC Photometer.
31BMS0001Photometer TWC10/30/201000:00:0011/01/201011:27:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201, TWC.
32BMSPhotometer TWC12/09/200912:11:0012/09/200914:22:00Due to a power outage at NREL, followed by bad power, the SRRL tracker UPS shutdown. The following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, NIP-RG780, LI201, TWC500, CUV4; and the following instruments were unshaded: 8-48, CM22, PSP, CUV4, PIR, CG4.
33BMS0001Photometer TWC07/30/200912:45:0007/30/200923:59:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201.
34BMSPhotometer TWC11/15/200814:02:0011/15/200814:32:00Due to a planned power outage at SRRL, the following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, LI201, TWC500, CUV4.
35BMS0001Photometer TWC09/30/200814:18:0010/01/200815:00:00The TWC photometer was not properly aligned with the sun after being relocated to the Sci-Tec tracker. Shims were added on 10/1/08 to improve alignment.
36BMSPhotometer TWC12/22/200708:41:0012/22/200711:24:00Due to a scheduled SRRL/Mesa Top power outage on 12/22/2007, the TSP-1, TSP-700, All Global UVA and UVB, and the TWC Photometer were without power, therefore data is either bad or bogus.
37BMS0001Photometer TWC05/25/200610:50:0005/25/200611:26:00Sun Follower III tracker was down due to power shutdown for installation of new SRRL equipment lift. The following instruments were NOT pointing at the SUN: NIP #1, NIP RG780, CUVB2, LI-201, TWC (500nm).
38BMSPhotometer TWC01/08/200400:00:0012/21/200514:53:00TWC Photometer was down due to detector failure. Detector and filter was replaced and the unit was returned to service. A new Vo calibration constant will need to be generated on the next clear day.
39BMSPhotometer TWC12/13/200007:21:0012/15/200009:46:00The 500nm Photometer was disconnected for repair. The instrument had been giving noisy signal at night between midnight and sunrise, and high voltage output (up to 13.5 V) at approx. 7:30 MST (almost every day), which caused the datalogger to put out missing values (-99999) due to over-ranging. This was caused by broken or loose connections in the instrument cable and depending on the position of the tracker the wires supplying power to the instrument shorted with the wires carrying the signal from the instrument. This wire has been repaired are re-oriented to help prevent future damage.
40BMSPhotometer TWC11/01/200008/23/200112:00:00An incorrect calibration factor of 0.487 W/mē/nm/mV was in the datalogger during this time period. The correct factor should have been 0.388 W/mē/nm/mV. This data was repaired on 6-14-2002, a correction is NOT required.